# A machine-learning photometric classifier for massive stars in nearby galaxies

### v1

@@ 4 Discussion
In this section, we compare our predictions with literature data to understand the performance of the classifier, explore the trends of the various classes with metallicity, and discuss the luminous RSGs and dusty YSGs in our sample.
4.1 Performance: comparison to literature
@@
Figure 3: Success rate vs. distance (top panels) and metallicity (bottom panels), using a uniform prior (left panels) and a unimodal beta distribution (right panels) with a peak corresponding to $77\pm 7$% (based on the performance of the classifier during developing). We notice a small decrease of the success rate with distance and a relatively flat behavior with metallicity, especially in the case where a prior is implemented (see Section 4.1 for details). The number of available classified sources from the literature is indicated by the size of the points and the colorbar on the right.
@@
To properly assess the performance of our classifier when comparing it with literature results (which are taken as ground truth) we considered two factors. The first was to provide a comparison by selecting the best candidates as defined from the quality criteria in Section 2.4 (i.e. final probability of ¿ 0.66 and band completeness ¿ 0.6). The second was to estimate the classification error, based on the probability distribution for each source.
Since the classifier provides a probability for each of the classes, the probability of two (or more) classes may be similar. In that case, although a single probability (and therefore class) is returned by the classifier, it does not mean that another class(es) is not possible (e.g. consider the case where we have 0.45 and 0.40 for a RSG and YSG class, respectively). To account for this we decided to estimate the class error by keeping all possible classes for which the probability did not differ significantly. For this, we first sorted all classes based on their probability (from highest to lowest). The first class in this sorting was always the one to be returned. If the ratio of the probability of the second to the first class differed more than a factor of two, it was rejected, and the first class was the only result. If the ratio of probabilities is less than two, the second class was added to the list of possible classes. We continued this process up to the point where the next class was rejected. Thus, we created a list of possible classes. When comparing our predictions to literature classes we identified the following cases:
First, an exact match when the first class of the list was the same or equivalent to the literature class. That was the case for straightforward classifications such as those used to train the classifier (see Paper I, table 2 and Section 3.2). For example, an A5 Ib or B9 Ia star predicted as a ’BSG’ would be an exact match. Similarly, a K2/3 I star or a carbon star predicted as a ’RSG’. Although the last pair (RSG and carbon star) is not a physical exact match, the classifier has been trained on a feature scheme (color indexes) for which these two classes overlap. In other words, due to the lack of features and data, the classifier cannot separate these two classes. Another example is when a rather broad classification is provided by the literature, such as ”emission line star”, which can be consistent with any of the classes of ’LBV’, ’BeBR’, ’WR’, and ’BSG’ [^10] This information was cross-matched with the source papers to verify their stellar nature. For example, the classification of ’EmObj’ from Bonanos et al. (2024) refers to objects that although display emission lines they are of uncertain nature. Therefore, these cases were considered uncertain.. For our classification scheme, these are equivalent cases and therefore correct whenever present.
Second, a match within the error when any class for the list of possible results was similar or equivalent to the literature class. That is the case when, for example, for a literature class of A5 II we had a predicted list of ’LBV’, ’YSG’, ’BSG’, or ’WR’. Although the first result ’LBV’ (which was the one returned as the final class) did not match the literature, there was a class within the accepted ones (i.e. ’BSG’) that matched the literature. Therefore, our prediction was correct within its error.
Third, there were cases where the literature result was too uncertain to safely match any of our classes. Such an example is the ’Composite’ classification assigned to the source B17 in WLM (Britavskiy et al., 2015), which does not provide any further information on how these sources can be compared to our classes.
Fourth, cases where none of the predicted classes matched the literature one. These are obvious cases such as a ’late G’ star or a carbon star predicted as ’WR’.
Fifth, a no-match result was returned when there was no match within 1″  between the literature source and our catalog. These sources were excluded from further analysis.
@@
We note here that when performing this comparison of spectral classification and the prediction from our photometric classifier, variability may introduce an offset. Especially for cases of sources with significant variability (e.g. LBVs, RSGs) the epochs of spectroscopic and photometric observations (which are not concurrent in the optical and IR) may lead to a mismatch between the spectral type and the prediction. For larger populations though (such as RSGs) we expect statistically consistent results.
@@
In Appendix B we provide in detail the statistics of matched sources per galaxy. In Fig. 3 we plot the success rate (corresponding to the number of correct and correct within error) over the subsample of literature sources, which matched best candidates from our catalog with distance and metallicity of each galaxy. We have taken extra care to deal with the errors in these fractions for two reasons. First, the fractions are bounded (0 to 1) so their error cannot extend beyond these values. Standard approaches of error estimate on the fractions (such as Wald or Wilson methods) would yield symmetric errors, leading to values exceeding the bounds (e.g. close to 1). Secondly, the sample size is different in each galaxy. Based on the current sample size and the number of correct predictions we get only one result (of the possible combinations) for the success rate, which follows a binomial distribution. To overcome all these and to avoid any Gaussian approximation, we have followed a Bayesian approach (see Appendix C for a more technical overview). We estimate the posterior probability of the success rate based on the likelihood provided by the binomial distribution (given the sample size and the correct predictions in each case) and a prior. The prior corresponds to the knowledge we have regarding the distribution of probable success rates for our classifier. We considered two options for the prior, a uniform one (i.e. single success rate value), and a unimodal beta distribution (plots in Fig. 3 left and right correspondingly; only galaxies with literature data and good candidates are plotted). This option translates to a success rate that has a single peak and drops off towards both bounds. We already have data for this from Paper I, in which we have concluded that the accuracy was 0.83 for the M31 and M33 galaxies (on which the classifier was trained) and 0.70 for the test galaxies (i.e. IC 1613, WLM, and Sextans A). We can therefore construct a distribution with a mode at a mean value of 0.77 and a variance of 0.01111111The difference between the two sample values is $\sim 0.07\%$, which corresponds to a variance of 0.005. As this value is indicative of how the distribution drops from the peak value we opted to relax it, and double the value to 0.01, to allow for a smoother distribution to include more probable success rates.. Given the likelihood and the prior we can construct the posterior probability distribution per galaxy. From this, we can extract basic statistical properties including asymmetric errors corresponding to the highest posterior density interval set at 95%.
@@
By examining the left panels of Fig. 3 we first notice that the success rate of the classifier drops with distance, although not significantly. This is expected, as the farther the galaxy is, the greater the level of confusion. The most extreme example is M83 for which we have,additionally, a small number of good candidates (fulfilling the criteria of Section 2.4). Interestingly, we also note that the classifier seems to be rather robust with metallicity. We achieve good results even at the lowest metallicity environments. There is an exception of three galaxies (namely NGC 7793, NGC 2403, and NGC 253) for which the success rate is lower than 0.4 (when using the uniform prior). These are the galaxies with the smallest number of classified samples for which we got very low success rates (e.g. one correct prediction out of the 6 sources in NGC 253, 0 correct out of 2 in NGC 7793, and 2 out of 7 in NGC 2403). These cases are also sensitive to the low-number statistics since a small change can lead to a significant change in the success rate. For example, in NGC 253 there were a couple of sources which were found uncertain (such as ’cluster’ or ’nebula’, which the classifier cannot predict). If we were to remove these two from the sample, then the original success rate would change from one out of 6 to one out of 4 sources (from 0.16 to 0.25), which is a significant change. On the other hand, Sextans B has only two classified sources that are both predicted correctly. Hence we ended up with a 100% success rate which was just a random realization of the binomial distribution.
@@
By incorporating the prior from the unimodal beta function that peaks at a mean success rate (as defined previously) we notice (see right panel of Fig. 3) that the values are updated and both extreme values (i.e. NGC 7793, NGC 253, NGC 2403, and Sextans A) are driven towards to more probable (and realistic) values, within the 60–94% success rate. The small decrease of the success rate with distance is still evident, but the performance of the classifier with respect to the metalliciy flattens. M31 and M33 stand out in these plots because they both have the largest numbers of sources by far ($\sim 430$ and $\sim 380$, respectively, resulting in small errors), and the best success rates. The latter is to be expected since the classifier has been trained on samples from these two galaxies.
@@
In summary, we conclude that our classifier performs exceptionally well even at lower metallicities (e.g. $0.07\,Z_{\odot}$ for Sextans A and Phoenix Dwarf), despite not being explicitly trained for them. There seems to be a greater dependence on distance (due to confusion) which makes it more efficient at distances smaller than 1.5 Mpc but with only a minor loss beyond 3 Mpc (where the furthest galaxies are located).
@@ 4.2 Populations with metallicity
Figure 4: The fraction of predicted population with metallicity per class. Despite the presence of large errors and important physical and observational biases, there are noticeable trends of each population with metallicity (see Section 4.2 for details). The symbol size and color for each galaxy reflects the corresponding sample size. For clarity we assign each galaxy to an integer ID, shown in the legend at the top.
@@
In this section, we examine the trend of each class separately with metallicity. But before presenting our results, we describe in detail our biases.
@@
We repeat here that our source selection was based on Spitzer point-source catalogs. That means that the catalog comprises of sources visible in the IR, which means that a certain fraction of some populations (e.g. OB main-sequence stars, stellar sources without dusty environments) are not included. Given this, we understand that the BSG class, for example, does not reflect the complete sample for any galaxy. That is evident if we compare the most recent catalog of OB stars in Sextans A by Lorenzo et al. (2022) with ours. We find matches for $\sim 28\%$ (of 106 sources in total, within a 1″  search radius). Another factor, influencing the fraction of populations we can observe, is the galaxy inclination. For example, NGC 55 is almost edge-on which means that a (significant possibly) fraction is not visible. While in other cases, such as in Sextans A which is a smaller and less crowded galaxy, we can pinpoint all sources. The difference between these two galaxies highlights that the galaxy’s type and star formation history have an important impact also. A recent star-formation event will lead to the birth of new stars which would (most probably and depending on the metallicity) lead to increased numbers of WR stars. As time passes these populations (along with OB main sequence stars) will decrease, and in turn the populations of more evolved phases such as YSG and RSG will increase.
@@
Following the above physical reasons there are several observational biases. One main constraint is the limitation in the pointings of Spitzer and its coverage, which affects the completeness. In Maravelias et al. (2023) we noted that in NGC 55 there are four known LBVs and our approach was successful to recover two of them. The other two were not recovered due to their location, which lay outside the observed Spitzer fields. Therefore, an unknown fraction of all populations are affected. This is more critical in the rarest cases (such as for LBVs and BeBRs) while probably not significant for the most populous ones (such as RSGs, and BSGs). Our classifier is based on the presence of photometry in certain Spitzer and Pan-STARRS bands. Additionally, many sources do not have photometry in all these bands, and, consequently, the quality of their predictions did not pass our quality criteria. Although this compensates for not adding erroneous predictions, there is an important fraction of sources not considered in our analysis.
@@
Considering all the above caveats, Fig. 4 presents the fraction of predictions per class over the total sample of only the best candidates. Our results are provided using the same approach for error determination described in the previous section (using the 95% interval), but we are using a separate prior for each class. The peak and the variance of the distribution for each class are provided from the results we have obtained in Paper I (c.f. table 5) [^12] For the LBV class we used the fractions and corresponding variance derived from the Support Vector Machine methods, the only non-zero result.. We removed galaxies without any good matches (NGC 55, NGC 300, Phoenix Dwarf, NGC 1313, NGC 5253).
@@
Starting with LBVs we cannot draw secure conclusions regarding their trend with metallicity, given their small numbers and large errors. There is some indication that lower-Z environments may host more LBVs (e.g. Sextans A).
@@
The B[e] supergiants (BeBR) is another rare class of objects, for which there are not many confirmed cases in low metallicity galaxies (see e.g. in Maravelias et al. 2023). These sources are (by definition) very bright in IR due to the dust formation in their complex circumstellar rings, which means that we should be able to recover almost their total populations across all galaxies. Therefore, the slight increase with decreasing metallicity should be real. Whether or not we expect a larger number of such objects in low metallicity environments is currently unknown (because of the lack of their formation channels; see e.g. Kraus 2019).
@@
Metallicity has a significant impact on stellar winds of hot stars, which are mainly driven by iron. Therefore, decreasing metallicity leads to less efficient winds, resulting in fewer stars managing to expel their outer layers to become WR. This is depicted in the plot by the decrease of the WR fraction with metallicity, which is slightly more evident below 0.4 Z. Given that the fraction of false positives within WR is quite high it is possible that their real fraction is even smaller than what depicted in the plot. Moreover, as they correspond to young massive stars their populations correlate also with regions of recent star formation events (such as starburst NGC 4214; Williams et al. 2011).
@@
As the metallicity decreases, stars lose their angular momentum much slower, preserving in this way their initial rotational speeds longer and increasing the internal mixing (the stars become more compact and have higher temperature and density in their centers). Consequently, there seems to be an extension of the main-sequence phase (hydrogen burning at the core; Georgy et al. 2013; Ekström et al. 2012). Simultaneously, it is possible to have He burning happening much earlier (i.e. higher temperatures), before the RSG phase (Yoon et al., 2008). Those stars will spend more time as YSGs rather than RSGs. Moreover, the more massive and luminous RSGs have strong winds that lead to a fast stripping of their envelopes, resulting in post-RSGs looking like BSG or YSG (Massey et al., 2023; Zapartas et al., 2024). Ultimately, these would lead to more BSGs and YSGs present at any time, than their corresponding numbers in higher metallicity galaxies. This is indeed observed in Fig. 4.
@@
The RSG population is not affected significantly (if any) by the metallicity (e.g. Antoniadis et al. 2024, 2025). Therefore, we would expect a rather flat trend or slight drop with metallicity. However, this is not what we see in the RSG panel of Fig. 4. We attribute this to an observational bias, as we constructed our catalog based on Spitzer sources corresponding mainly to evolved and dusty sources. At lower metallicity galaxies, more stars spend their majority of their time in the BSG/YSG phase, which means that the number of these sources with dust will drop. Consequently, the ratios of RGSs over the total number of IR sources (which includes fewer actually observed BSG/YSG as the metallicity decreases) lead in an upward trend.

### v2

@@ 4 Discussion
In this section, we compare our predictions with the literature data to understand the performance of the classifier. We also explore the trends of the various classes with metallicity, and discuss the luminous RSGs and dusty YSGs in our sample.
4.1 Performance: comparison to the literature
@@
Figure 3: Success rate vs. distance (top panels) and metallicity (bottom panels), using a uniform prior (left panels) and a unimodal beta distribution (right panels) with a peak corresponding to $77\pm 7$% (based on the performance of the classifier during development). We notice a small decrease in the success rate with distance and a relatively flat behavior with metallicity, especially in the case where a prior is implemented (see Sect. 4.1 for details). The number of available classified sources from the literature is indicated by the size of the points and the colorbar on the right.
@@
To properly assess the performance of our classifier when comparing it with the literature results (which are taken as ground truth), we considered two factors. The first was to provide a comparison by selecting the best candidates as defined from the quality criteria in Sect. 2.4 (i.e., the final probability of ¿ 0.66 and band completeness ¿ 0.6). The second was to take into account the classification uncertainty by combining the full information from the probability distribution across all classes.
Since the classifier provides a probability for each of the classes, the probability of two (or more) classes may be similar. In that case, although a single final class is returned as the best prediction, it does not exclude others (e.g., probabilities of 0.45 and 0.40 for a RSG and YSG class, respectively). To account for this, we employed the categorical cross entropy (CCE) metric. It estimates the entropy contribution of each class based on the probability for the particular class returned from the classifier. This is multiplied with the one-hot encoder of the literature spectral type, which transforms a category to a vector with a length equal to the number of available classes. Only the class corresponding to the literature actually acts on the probability list. In other words, for each source:
$CCE=-\sum_{k=1}^{N}[y_{\textrm{{true}}}\times\ln(y_{\textrm{{pred}}})],$
where $y_{\textrm{pred}}$ is the array of the probability returned by the classifier for each class, and $y_{\textrm{true}}$ is the vector that maps the literature type to the prediction list, both of them with dimensions of $1\times N$ (in our case $N=7$ classes). For example, if our list of predictions has the following order:
$y_{\textrm{{true}}}=[p_{\textrm{BSG}},p_{\textrm{BeBR}},p_{\textrm{GAL}},p_{\textrm{LBV}},p_{\textrm{RSG}},p_{\textrm{WR}},p_{\textrm{YSG}}].$
If the literature provides a RSG classification, the one-hot encoded vector is
$y_{true}=[0,0,0,0,1,0,0].$
Then, the CCE considers only the probability for the RSG class, and reduces to
$CCE=-[1\times\ln(p_{\textrm{RSG}})].$
In cases where the literature classification was broader (e.g., a ”blue” or ”hot” star that is compatible with any of the predicted classes of BSG, LBV, WR, or BeBR), we used a soft one-hot encoding where the probability is equally split across all these classes, so that
$CCE=-[0.25\times\ln(p_{\textrm{BSG}})+0.25\times\ln(p_{\textrm{BeBR}})$
$+0.25\times\ln(p_{\textrm{LBV}})+0.25\times\ln(p_{\textrm{WR}})].$
By construction, lower CCE values indicate more secure classifications (the best being 0). However, this is far from true in practice. For this, we needed to define the threshold below which we considered our predictions secure. We already knew that our cutoff on probability was $p=0.66$. At the same time, such a probability actually meant that the prediction of the classifier was a single class with a relatively high probability (with probability at least double the probability of the next predicted class [^9] In the case of two predicted classes, if the first (returned) class has a probability twice as large than the second one, we will have: $p_{1}+p_{2}=1\Rightarrow 2\times p_{2}+p_{1}=1\Rightarrow p_{2}=0.33$, so that $p1=0.66$.). On the other hand, a random classifier with seven classes returns a probability of $1/7\approx 0.14$. Therefore, we defined three regions that represented confidence in our predictions: a. ”secure”: $p>0.66\Rightarrow CCE<0.42$, b. ”candidate”: $0.14<p\leq 0.66\Rightarrow 0.42\geq CCE>1.95$, and c. ”uncertain”: $p\leq 0.14\Rightarrow CCE\geq 1.95$. In that way, our estimate of robustness accounted for the contribution from all classes, in conjunction with the information from the literature.
We used the above approach in all sources for which a positional match between our catalog and the literature data was found within 1″. Sources without a match were excluded. Sources with ambiguous or uncertain literature classifications (such as the ”composite” classification assigned to source B17 in WLM Britavskiy et al. 2015) were also excluded from further consideration, as they lack a reliable ground truth for validation. However, we retained these sources in the published catalog for completeness, but did not use them to assess classification performance.
@@
We note here that when this comparison of spectral classification and the prediction from our photometric classifier is performed, variability may introduce an offset. Especially in cases of sources with significant variability (e.g., LBVs, RSGs), the epochs of spectroscopic and photometric observations (which are not concurrent in the optical and IR) may lead to a mismatch between the spectral type and the prediction. However, for larger populations (such as RSGs) we expect statistically consistent results.
@@
In Appendix C we provide in detail the statistics of matched sources per galaxy. In Fig. 3 we plot the success rate (corresponding to the number of secure and candidate classifications) over the subsample of literature sources, which matched the best candidates from our catalog with the distance and metallicity of each galaxy. We took extra care to deal with the errors in these fractions for two reasons. First, the fractions are bounded (0 to 1), so their error cannot extend beyond these values. Standard approaches of error estimate on the fractions (such as Wald or Wilson methods) yield symmetric errors, leading to values exceeding the bounds (e.g., close to 1). Second, the sample size is different in each galaxy. Based on the current sample size and the number of correct predictions (i.e., the total number of secure and candidate classifications, as defined previously), we get only one result (of the possible combinations) for the success rate, which follows a binomial distribution. To overcome all these and to avoid any Gaussian approximation, we followed a Bayesian approach (see Appendix D for a more technical overview). We estimated the posterior probability of the success rate based on the likelihood provided by the binomial distribution (given the sample size and the correct predictions in each case) and a prior. The prior corresponds to our knowledge of the distribution of probable success rates for our classifier. We considered two options for the prior, a uniform one (i.e., single success rate value) and a unimodal beta distribution (plots in Fig. 3 left and right correspondingly; only galaxies with literature data and good candidates, i.e., fulfilling the criteria of Sect. 2.4, were plotted). This option translates to a success rate that has a single peak and drops off toward both bounds. We already had data for this from Paper I, in which we concluded that the accuracy was 0.83 for the M31 and M33 galaxies (on which the classifier was trained) and 0.70 for the test galaxies (i.e., IC 1613, WLM, and Sextans A). We therefore constructed a distribution with a mode at a mean value of 0.77 and a variance of 0.01101010The difference between the two sample values is $\sim 0.07\%$, which corresponds to a variance of 0.005. As this value is indicative of how the distribution drops from the peak value we opted to relax it, and double the value to 0.01, to allow for a smoother distribution to include more probable success rates.. Given the likelihood and the prior, we constructed the posterior probability distribution per galaxy. From this we extracted basic statistical properties including asymmetric errors corresponding to the highest posterior density interval set at 95%.
@@
By examining the left panels of Fig. 3 we first notice that the success rate of the classifier drops with distance, although not significantly. This is expected, as the farther the galaxy is, the greater the level of confusion. The most extreme example is M83 for which we have, additionally, a small number of good candidates, but with relatively broader classifications from the literature (mainly Hii regions and emission stars). Interestingly, we also note that the classifier seems to be rather robust with metallicity. We achieve good results even at the lowest metallicity environments. There is an exception of three galaxies (namely NGC 7793, NGC 2403, and NGC 253) for which the success rate is lower than 0.4 (when using the uniform prior). These are the galaxies with the smallest number of classified samples for which we get very low success rates (e.g., one secure prediction out of the four sources in NGC 253, zero correct out of two in NGC 7793, and two secure and three candidates out of seven in NGC 2403). These cases are also sensitive to the low-number statistics, since a small change can lead to a significant change in the success rate. On the other hand, Sextans B has only two classified, of which we correctly matched and predicted only one. Hence, we ended up with a 100% success rate, which was just a random realization of the binomial distribution.
@@
By incorporating the prior from the unimodal beta function that peaks at a mean success rate (as defined previously) we notice (see the right panel of Fig. 3) that the values were updated and both extreme values (i.e., NGC 7793, NGC 253, NGC 2403, and Sextans B) are driven toward more probable (and realistic) values, within the 60–94% success rate. The small decrease in the success rate with distance was still evident, but the performance of the classifier with respect to the metallicity flattened. M31 and M33 stand out in these plots because they both have the highest number of sources by far ($\sim 430$ and $\sim 380$, respectively, resulting in small errors) and the best success rates. The latter is expected since the classifier has been trained on samples from these two galaxies.
@@
In summary, we conclude that our classifier performs exceptionally well even at lower metallicities (e.g., $0.07\,Z_{\odot}$ for Sextans A and Phoenix Dwarf), despite not being explicitly trained for them. There seems to be a greater dependence on distance (due to confusion), which makes it more efficient at distances smaller than 1.5 Mpc but with only a minor loss beyond 3 Mpc (where the furthest galaxies are located).
@@ 4.2 Populations with metallicity
Figure 4: Fraction of predicted population with metallicity per class. Despite the presence of large errors and important physical and observational biases, there are noticeable trends of each population with metallicity (see Sect. 4.2 for details). The symbol size and color for each galaxy reflects the corresponding sample size. For clarity we assign each galaxy to an integer ID, shown in the legend at the top.
@@
In this section, we examine the trend of each class separately with metallicity. However, before presenting our results, we describe our biases in detail.
@@
We repeat here that our source selection was based on Spitzer point source catalogs. That means that the catalog comprises of sources visible in the IR, which means that a certain fraction of some populations (e.g., OB main-sequence stars, stellar sources without dusty environments) are not included. Given this, we understand that the BSG class, for example, does not reflect the complete sample for any galaxy. This is evident if we compare the most recent catalog of OB stars in Sextans A by Lorenzo et al. (2022) with ours. We find matches for $\sim 28\%$ (of 106 sources in total, within a 1″  search radius). Another factor influencing the fraction of populations that we can observe is the galaxy inclination. For example, NGC 55 is almost edge-on, which means that a (possibly significant) fraction is not visible. In other cases, such as in Sextans A, which is a smaller and less crowded galaxy, we can pinpoint all sources. The difference between these two galaxies highlights that the galaxy’s type and star formation history also have an important impact. A recent star-formation event will lead to the birth of new stars, which will probably (depending on the metallicity) lead to an increased number of WR stars. As time passes these populations (along with OB main sequence stars) will decrease, and in turn the populations of more evolved phases such as YSG and RSG will increase.
@@
Following the above physical reasons, there are several observational biases. One main constraint is the limitation in the pointings of Spitzer and its coverage, which affects the completeness. In Maravelias et al. (2023) we noted that in NGC 55 there are four known LBVs and our approach was successful in recovering two of them. The other two were not recovered due to their location, which lay outside the observed Spitzer fields. Therefore, an unknown fraction of all populations are affected. This is more critical in the rarest cases (such as for LBVs and BeBRs) while probably not significant for the most populous ones (such as RSGs and BSGs). Our classifier is based on the presence of photometry in certain Spitzer and Pan-STARRS bands. Additionally, many sources do not have photometry in all these bands, and consequently, the quality of their predictions did not pass our quality criteria. Although this compensates for not adding erroneous predictions, there is an important fraction of sources not considered in our analysis.
@@
Considering all the above caveats, Fig. 4 presents the fraction of predictions per class over the total sample of only the best candidates. Our results are provided using the same approach for error determination described in the previous section (using the 95% interval), but we use a separate prior for each class. The prior is derived from the unimodal beta function of each class, which peaks at the mean success rate and with a variance according to the results obtained in Paper I (cf., Table 5) [^11] For the LBV class we used the fractions and corresponding variance derived from the support vector machine methods, the only nonzero result.. We removed galaxies without any good matches (NGC 55, NGC 300, Phoenix Dwarf, NGC 1313, and NGC 5253).
@@
Starting with LBVs, we cannot draw secure conclusions regarding their trend with metallicity given their small numbers and large errors. There is though indication that lower-Z environments may host more LBVs (e.g., Sextans A).
@@
The B[e] supergiants (BeBR) is another rare class of objects, for which there are not many confirmed cases in low metallicity galaxies (see e.g., in Maravelias et al. 2023). These sources are, by definition, very bright in IR due to the dust formation in their complex circumstellar rings, which means that we should be able to recover almost their total populations across all galaxies. Therefore, the slight increase with decreasing metallicity should be real. Whether or not we expect a higher number of such objects in low metallicity environments is currently unknown (because of the lack of their formation channels; see e.g., Kraus 2019).
@@
Metallicity has a significant impact on stellar winds of hot stars, which are mainly driven by iron. Therefore, decreasing metallicity leads to less efficient winds, resulting in fewer stars managing to expel their outer layers to become WR. This is depicted in the plot by the decrease in the WR fraction with metallicity, which is slightly more evident below 0.4 Z. Given that the fraction of false positives within WR is quite high, it is possible that their real fraction is even smaller than that depicted in the plot. Moreover, as they correspond to young massive stars, their populations also correlate with regions of recent star formation events (such as starburst NGC 4214; Williams et al. 2011).
@@
As the metallicity decreases, stars lose their angular momentum much slower, preserving in this way their initial rotational speeds longer and increasing the internal mixing (the stars become more compact and have higher temperature and density in their centers). Consequently, there seems to be an extension of the main-sequence phase (hydrogen burning at the core; Georgy et al. 2013; Ekström et al. 2012). Simultaneously, it is possible for He burning to occur much earlier (i.e., higher temperatures), before the RSG phase (Yoon et al., 2008). Those stars will spend more time as YSGs than RSGs. Moreover, the more massive and luminous RSGs have strong winds that lead to a fast stripping of their envelopes, resulting in post-RSGs that resemble BSGs or YSGs (Massey et al., 2023; Zapartas et al., 2025). Ultimately, these lead to more BSGs and YSGs present at any time than their corresponding numbers in higher metallicity galaxies. This is, indeed, observed in Fig. 4.
@@
The RSG population is not affected significantly (if at all) by metallicity (e.g., Antoniadis et al. 2024, 2025). Therefore, we expect a rather flat trend or a slight drop in metallicity. However, this is not what we see in the RSG panel of Fig. 4. We attribute this to observational bias, as we constructed our catalog based on Spitzer sources corresponding mainly to evolved and dusty sources. At lower metallicity galaxies, more stars spend their majority of the time in the BSG or YSG phase, which means that the number of these sources with dust will drop. Consequently, the ratios of RGSs over the total number of IR sources (which include fewer actually observed BSGs or YSGs as the metallicity decreases) lead to an upward trend.